Calculate the Opportunity Cost (O/C) of completing ECON1 successfully for WINTER 2019?

This Opportunity cost should include the obvious cost (E.g. Book, Tuition, parking etc) and hidden cost (value of your time: how many hours per week * 12 weeks: What you will be doing and what it will be worth to you. E.g.

Hidden cost is the O/C of time: what kind of uses of your time, e.g. let's say you spend 100 hours to finish the Econ 1 class successfully.

What else could you have done in these 100 hours and how much was that worth to you. Again, let's say 20 hrs spent with your family. How much is each hour worth $100. So 20 * $100 = $2000 and so on. It could be 25 hours of work, multiply with the wage rate etc.

PreviousNext
Answer #1

The opportunity cost of completing the Econ1 course for winter 2019 progam can be divided into the explicit cost and implicit cost.

Explicit costs are as follows:

Tuition fee = $5000/-

Cost of Books = $500/-

Cost of food & lodging = $1000 per month = 1000*3 = $3000

Cost of conveyance, parking and other petty expenses = $500 per month = 500*3 = $1500

So, total explicit cost = 5000+500+3000+1500 = $10000

Implicit costs (hidden cost) are as follows:

'cost of per hour of work = $20

Loss of hours of work = 20 hours per week

Cost of loss of work per week = 20*20 = $400

So, total cost of loss of work = 400*12 = $4800

Cost of attending one conference = $100

Total conference to be attended = 5

So, total cost of attending the conference = 100*5 = $500

Then, total implicit cost = 4800+500 = $5300

Hence,

Total opportunity cost = 10000 + 5300 = $15300